# Continuous Product Graph Neural Networks

### v1

@@ Abstract
Processing multidomain data defined on multiple graphs holds significant potential in various practical applications in computer science. However, current methods are mostly limited to discrete graph filtering operations. Tensorial partial differential equations on graphs (TPDEGs) provide a principled framework for modeling structured data across multiple interacting graphs, addressing the limitations of the existing discrete methodologies. In this paper, we introduce Continuous Product Graph Neural Networks (CITRUS) that emerge as a natural solution to the TPDEG. CITRUS leverages the separability of continuous heat kernels from Cartesian graph products to efficiently implement graph spectral decomposition. We conduct thorough theoretical analyses of the stability and over-smoothing properties of CITRUS in response to domain-specific graph perturbations and graph spectra effects on the performance. We evaluate CITRUS on well-known traffic and weather spatiotemporal forecasting datasets, demonstrating superior performance over existing approaches.
Keywords Graph Neural Networks  $\cdot$ Partial Differential Equations  $\cdot$ Product Graphs.
@@ 1 Introduction
The main challenge for learning from multidomain graph data is creating efficient frameworks that model joint interactions across domain-specific graphs. Previous work in this area has utilized discrete graph filtering operations in product graphs (PGs) [11, 10] from the field of graph signal processing (GSP) [12]. However, these methods inherit the well-known issues of over-smoothing and over-squashing from regular graph neural networks (GNNs) [13, 14, 15], which restricts the graph’s receptive field and hinders long-range interactions [16]. Additionally, these methods often require computationally intensive grid searches to tune hyperparameters and are typically limited to two-domain graph data, such as spatial and temporal dimensions [10, 17, 11].
@@ 2 Related Work
Fundamentals. The study of PGs is a well-established field in mathematics and signal processing [21, 22]. However, to the best of our knowledge, the graph-time convolutional neural network (GTCNN) [10, 17] model is the only GNN-based framework addressing neural networks in PGs. GTCNN defines a discrete filtering operation in PGs to jointly process spatial and temporal data for forecasting applications. Due to the discrete nature of the graph polynomial filters in GTCNN, an expensive grid search is required to correctly tune the graph filter orders, leading to significant computational overhead. Additionally, this discrete aspect restricts node-wise receptive fields, making long-range communication challenging [16, 18]. Furthermore, GTCNN is limited to two-factor graphs (space and time), limiting its applicability to general multidomain graph data. In contrast, CITRUS overcomes these limitations by i) employing continuous graph filtering, and ii) providing a general framework for any number of factor graphs.
Applicability. We explore spatiotemporal analysis as a specific application of PG processing. Early works in GSP for spatiotemporal forecasting, such as the graph polynomial vector auto-regressive moving average (GP-VARMA) and GP-VAR models [23], do not utilize PGs. Modern GNN-based architectures can be broadly classified into TTS and STT frameworks. TTS networks, due to their sequential nature, often encounter propagation information bottlenecks during training [8]. Conversely, STT architectures enhance node representations first and then integrate temporal information, but the use of recurrent neural networks (RNNs) in the aggregation stages leads to high computational complexity. Additionally, there are T&S frameworks designed for simultaneous learning of spatiotemporal representations [8, 24], where the fusion of temporal and spatial modules is crucial. However, similar to STT networks, T&S frameworks also face challenges with computational complexity [8] and their discretized, block-wise nature, which can limit adaptive graph receptive fields [18, 24].
CITRUS leverages joint multidomain learning by exploiting continuous separable heat kernels as factor-based functions defined on a Cartesian product graph. The continuity and differentiability of these filters allow the graph receptive fields to be learned adaptively during the training process, eliminating the need for a grid search. Additionally, CITRUS maintains low computational complexity by relying on the spectral decomposition of the small factor graphs [16, 20]. Furthermore, unlike non-graph approaches, the number of learnable parameters in CITRUS is independent of the factor graphs, ensuring scalability.
@@ 3 Methodology and Theoretical Analysis
Preliminaries and notation. An undirected and weighted graph $\mathcal{G}$ with $N$ vertices can be stated as $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathbf{A}\}$. $\mathcal{V}$ and $\mathcal{E}$ denote the sets of the vertices and edges, respectively. $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the adjacency matrix with $\{\mathbf{A}_{ij}=\mathbf{A}_{ji}\geq 0\}_{i,j=1}^{N}$ representing the connection between vertices, and $\{\mathbf{A}_{ii}=0\}_{i=1}^{N}$ states that there are no self-loops in $\mathcal{G}$. The graph Laplacian $\mathbf{L}\in\mathbb{R}^{N\times N}$ is defined as $\mathbf{L}=\mathbf{D}-\mathbf{A}$, where $\mathbf{D}=\operatorname{\operatorname{diag}}(\mathbf{A}\mathbf{1})$ is the diagonal degree matrix and $\mathbf{1}$ is the all-one vector. A graph signal is a function that maps a set of nodes to the real values $x:\mathcal{V}\to\mathbb{R}$, and therefore we can represent it as $\mathbf{x}=[x_{1},\ldots,x_{N}]^{\top}$. Using the Kronecker product $\otimes$, Kronecker sum (aka Cartesian product) $\oplus$, and the Laplacian factors $\{\mathbf{L}_{p}\in\mathbb{R}^{N_{p}\times N_{p}}\}$, we have the following definitions:
@@
where the Laplacian matrix of the Cartesian product between two-factor graphs can be stated as:
@@
with $\mathbf{I}_{n}$ the identity matrix of size $n$. See Figure 1(a) for an example of a Cartesian graph product between three factor graphs. A similar relationship to (2) also holds for adjacency matrices [25]. We define a $D$-dimensional tensor as $\underline{\mathbf{U}}\in\mathbb{R}^{N_{1}\times\ldots\times N_{D}}$. The matrix $\underline{\mathbf{U}}_{(i)}\in\mathbb{R}^{N_{i}\times\left[\prod_{r=1,r\neq i}^{D}{N_{r}}\right]}$ is the $i$-th mode matricization of $\underline{\mathbf{U}}$ [26]. The mode-$i$ tensorial multiplication of a tensor $\underline{\mathbf{U}}$ by a matrix $\mathbf{X}\in\mathbb{R}^{m\times N_{i}}$ is denoted by $\underline{\mathbf{G}}=\underline{\mathbf{U}}\times_{i}\mathbf{X}$, where $\underline{\mathbf{G}}\in\mathbb{R}^{N_{1}\times\ldots\times N_{i-1}\times m\times N_{i+1}\times\ldots\times N_{D}}$ [26]. For further information about the tensor matricization and mode-$n$ product, please refer to [26, 27] (especially Sections 2.4 and 2.5 in [26]). All the proofs of theorems, propositions, and lemmas of this paper are provided in the Appendix A.
@@ 3.2 Stability Analysis
Consider a Cartesian product graph (without isolated nodes) with the normalized true and perturbed Laplacians $\hat{\mathbf{L}}_{\diamond}$ and $\tilde{\hat{\mathbf{L}}}_{\diamond}$. Also, assume we have the perturbation model $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{E}$ with factor error bounds $\{{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\mathbf{E}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{p}\leq\varepsilon_{p}\}_{p=1}^{P}$. Then, the stability bound on the true and perturbed outputs of CITRUS , i.e., $\varphi(u,t)$ and $\tilde{\varphi}(u,t)$, respectively, can be described by the summation of the factor stability bounds as:
$\|\varphi(u,t)-\tilde{\varphi}(u,t)\|\leq\sum_{p=1}^{P}{\mathcal{O}(\varepsilon_{p})}.$ (11)
@@ 3.3 Over-smoothing Analysis
$\text{where}\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>E(\mathbf{U})\vcentcolon=\frac{1}{2}\sum_{(i,j)\in\mathcal{E}}{\left\|\frac{\mathbf{u}_{i}}{\sqrt{\text{deg}_{i}}}-\frac{\mathbf{u}_{j}}{\sqrt{\text{deg}_{j}}}\right\|^{2}}=\text{tr}(\mathbf{U}^{\top}\hat{\mathbf{L}}\mathbf{U}).$ (13)
Since (13) is based on the normalized Laplacians, we also need to express the product adjacency and Laplacian based on the normalized factor graphs.
@@
By constructing the normalized product adjacency as $\hat{\mathbf{A}}\vcentcolon=\frac{1}{P}\oplus_{p=1}^{P}{\hat{\mathbf{A}}_{p}}=\oplus_{p=1}^{P}{\left(\frac{\hat{\mathbf{A}}_{p}}{P}\right)}$, a similar Cartesian form for the normalized product Laplacian also holds as the following, where the spectrum is spanned across the interval of $[0,2]$:
$\hat{\mathbf{L}}\vcentcolon=\frac{1}{P}\oplus_{p=1}^{P}{\hat{\mathbf{L}}_{p}}=\oplus_{p=1}^{P}{\left(\frac{\hat{\mathbf{L}}_{p}}{P}\right)}.$ (14)
Lemma 3.8 shows that the Cartesian product graph of the normalized Laplacians (with a normalizing scalar $P$) leads to a valid normalized Laplacian. Next, by considering the product structure w.r.t. the underlying normalized Laplacians and also the tensorial structure of $\underline{\mathbf{U}}_{t}\in\mathbb{R}^{N_{1}\times N_{2}\times\ldots\times N_{P}\times F}$, the Dirichlet energy, which is a normalized version of the smoothness term in the GSP literature [12, 31], can be stated as [32, 25, 33]:
$E(\underline{\mathbf{U}}_{t})=\frac{1}{P}\sum_{f=1}^{F}{\sum_{p=1}^{P}{\text{tr}({\tilde{\mathbf{U}}^{\top}_{f(p)}}{\hat{\mathbf{L}}_{p}}{\tilde{\mathbf{U}}_{f(p)}})}},$ (15)
where ${\tilde{\mathbf{U}}_{f(p)}}=\underline{\mathbf{U}}_{t}[:,\ldots,:,f]_{(p)}\in\mathbb{R}^{N_{p}\times\prod_{i\neq p}^{P}{N_{i}}}$, i.e., the $p$-th mode matricization of the $f$-th slice of $\underline{\mathbf{U}}_{t}$ in its $(P+1)$-th dimension. We observe in (15) that the kernel separability of Cartesian heat kernels also leads to the separability of analyzing over-smoothing in different modes of the data at hand. This is useful in cases of facing factor graphs with different characteristics. Next, we formally analyze over-smoothing in CITRUS. In our case, for the $l$-th layer with $H_{l}$ hidden MLP layers, non-linear activations $\sigma(\cdot)$ and learnable weight matrices $\{\mathbf{W}_{lh}\}_{h=1}^{H_{l}}$, we define:
@@
For the product graph case $\hat{\mathbf{L}}=\frac{1}{P}\oplus_{p=1}^{P}{\hat{\mathbf{L}}_{p}}$ with domain-specific receptive field $t^{(i)}$ for the $i$-th factor graph and the activations $\sigma(\cdot)$ in (16) being ReLU or Leaky ReLU, we have:
@@
Theorem 3.9 shows that the factor graph with the smallest non-zero eigenvalue (spectral gap) multiplied by its receptive field dominates the overall over-smoothing. The less the spectral gap, the less probability the factor graph is connected [13]. So, we can focus on this specific factor graph instead of the resulting massive product graph. These theoretical findings are experimentally validated in Section 4.2.
@@ 4 Experimental Results
In this section, we experimentally validate the theoretical discussions regarding the stability and over-smoothing analysis, and use CITRUS in real-world spatiotemporal forecasting tasks. Similarly, we conduct ablation studies regarding CITRUS in spatiotemporal forecasting. We compare CITRUS against several state-of-the-art methods in forecasting. The results on the case of more than two factor graphs and descriptions of the state-of-the-art are provided in Appendix B and C, respectively.

### v2

@@ Abstract
Processing multidomain data defined on multiple graphs holds significant potential in various practical applications in computer science. However, current methods are mostly limited to discrete graph filtering operations. Tensorial partial differential equations on graphs (TPDEGs) provide a principled framework for modeling structured data across multiple interacting graphs, addressing the limitations of the existing discrete methodologies. In this paper, we introduce Continuous Product Graph Neural Networks (CITRUS) that emerge as a natural solution to the TPDEG. CITRUS leverages the separability of continuous heat kernels from Cartesian graph products to efficiently implement graph spectral decomposition. We conduct thorough theoretical analyses of the stability and over-smoothing properties of CITRUS in response to domain-specific graph perturbations and graph spectra effects on the performance. We evaluate CITRUS on well-known traffic and weather spatiotemporal forecasting datasets, demonstrating superior performance over existing approaches. The implementation codes are available at https://github.com/ArefEinizade2/CITRUS.
@@ 1 Introduction
The main challenge for learning from multidomain graph data is creating efficient frameworks that model joint interactions across domain-specific graphs [10, 11]. Previous work in this area has utilized discrete graph filtering operations in product graphs (PGs) [12, 10] from the field of graph signal processing (GSP) [13]. However, these methods inherit the well-known issues of over-smoothing and over-squashing from regular graph neural networks (GNNs) [14, 15, 16], which restricts the graph’s receptive field and hinders long-range interactions [17]. Additionally, these methods often require computationally intensive grid searches to tune hyperparameters and are typically limited to two-domain graph data, such as spatial and temporal dimensions [10, 18, 12, 19].
@@ 2 Related Work
Fundamentals. The study of PGs is a well-established field in mathematics and signal processing [23, 24]. However, to the best of our knowledge, the graph-time convolutional neural network (GTCNN) [10, 18] model is the only GNN-based framework addressing neural networks in PGs. GTCNN defines a discrete filtering operation in PGs to jointly process spatial and temporal data for forecasting applications. Due to the discrete nature of the graph polynomial filters in GTCNN, an expensive grid search is required to correctly tune the graph filter orders, leading to significant computational overhead. Additionally, this discrete aspect restricts node-wise receptive fields, making long-range communication challenging [17, 20]. Furthermore, GTCNN is limited to two factor graphs (space and time), limiting its applicability to general multidomain graph data. In contrast, CITRUS overcomes these limitations by i) employing continuous graph filtering, and ii) providing a general framework for any number of factor graphs.
Applicability. We explore spatiotemporal analysis as a specific application of PG processing. Early works in GSP for spatiotemporal forecasting, such as the graph polynomial vector auto-regressive moving average (GP-VARMA) and GP-VAR models [25], do not utilize PGs. Modern GNN-based architectures can be broadly classified into TTS and STT frameworks. TTS networks, due to their sequential nature, often encounter propagation information bottlenecks during training [8]. Conversely, STT architectures enhance node representations first and then integrate temporal information, but the use of recurrent neural networks (RNNs) in the aggregation stages leads to high computational complexity. Additionally, there are T&S frameworks designed for simultaneous learning of spatiotemporal representations [8, 26], where the fusion of temporal and spatial modules is crucial. However, similar to STT networks, T&S frameworks also face challenges with computational complexity [8] and their discretized block-wise nature, which can limit adaptive graph receptive fields [20, 26].
CITRUS leverages joint multidomain learning by exploiting continuous separable heat kernels as factor-based functions defined on a Cartesian product graph. The continuity and differentiability of these filters allow the graph receptive fields to be learned adaptively during the training process, eliminating the need for a grid search. Additionally, CITRUS maintains low computational complexity by relying on the spectral decomposition of the factor graphs [17, 22]. Furthermore, unlike non-graph approaches, the number of learnable parameters in CITRUS is independent of the factor graphs, ensuring scalability.
@@ 3 Methodology and Theoretical Analysis
Preliminaries and notation. An undirected and weighted graph $\mathcal{G}$ with $N$ vertices can be stated as $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathbf{A}\}$. $\mathcal{V}$ and $\mathcal{E}$ denote the sets of the vertices and edges, respectively. $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the adjacency matrix with $\{\mathbf{A}_{ij}=\mathbf{A}_{ji}\geq 0\}_{i,j=1}^{N}$ representing the connection between vertices, and $\{\mathbf{A}_{ii}=0\}_{i=1}^{N}$ states that there are no self-loops in $\mathcal{G}$. The graph Laplacian $\mathbf{L}\in\mathbb{R}^{N\times N}$ is defined as $\mathbf{L}=\mathbf{D}-\mathbf{A}$, where $\mathbf{D}=\operatorname{\operatorname{diag}}(\mathbf{A}\mathbf{1})$ is the diagonal degree matrix and $\mathbf{1}$ is the all-one vector. A graph signal is a function that maps a set of nodes to the real values $x:\mathcal{V}\to\mathbb{R}$, and therefore we can represent it as $\mathbf{x}=[x_{1},\ldots,x_{N}]^{\top}$. For a vector $\mathbf{a}=[a_{1},\ldots,a_{N}]^{\top}\in\mathbb{R}^{N\times 1}$, $e^{\mathbf{a}}:=[e^{a_{1}},\ldots,e^{a_{N}}]^{\top}$. The vectorization operator is shown as $\operatorname{\operatorname{vec}}(.)$. Using the Kronecker product $\otimes$, Kronecker sum (aka Cartesian product) $\oplus$, and the Laplacian factors $\{\mathbf{L}_{p}\in\mathbb{R}^{N_{p}\times N_{p}}\}_{p=1}^{P}$, we have the following definitions:
@@
where the Laplacian matrix of the Cartesian product between two factor graphs can be stated as:
@@
with $\mathbf{I}_{n}$ the identity matrix of size $n$. See Figure 1(a) for an example of a Cartesian graph product between three factor graphs. A similar relationship to (2) also holds for adjacency matrices [24, 27]. We define a $D$-dimensional tensor as $\underline{\mathbf{U}}\in\mathbb{R}^{N_{1}\times\ldots\times N_{D}}$. The matrix $\underline{\mathbf{U}}_{(i)}\in\mathbb{R}^{N_{i}\times\left[\prod_{r=1,r\neq i}^{D}{N_{r}}\right]}$ is the $i$-th mode matricization of $\underline{\mathbf{U}}$ [28]. The mode-$i$ tensorial multiplication of a tensor $\underline{\mathbf{U}}$ by a matrix $\mathbf{X}\in\mathbb{R}^{m\times N_{i}}$ is denoted by $\underline{\mathbf{G}}=\underline{\mathbf{U}}\times_{i}\mathbf{X}$, where $\underline{\mathbf{G}}\in\mathbb{R}^{N_{1}\times\ldots\times N_{i-1}\times m\times N_{i+1}\times\ldots\times N_{D}}$ [28]. For further information about the tensor matricization and mode-$n$ product, please refer to [28, 29] (especially Sections 2.4 and 2.5 in [28]). All the proofs of theorems, propositions, and lemmas of this paper are provided in the Appendix A.
@@ 3.2 Stability Analysis
Consider a Cartesian product graph (without isolated nodes) with the normalized true and perturbed Laplacians $\hat{\mathbf{L}}_{\diamond}$ and $\tilde{\hat{\mathbf{L}}}_{\diamond}$ in (6). Also, assume we have the perturbation model $\tilde{\mathbf{A}}_{p}=\mathbf{A}_{p}+\mathbf{E}_{p}$ with factor error bounds $\{{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\mathbf{E}_{p}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}\leq\varepsilon_{p}\}_{p=1}^{P}$ in Proposition 3.6. Then, the stability bound on the true and perturbed outputs of CITRUS , i.e., $\varphi(u,t)$ and $\tilde{\varphi}(u,t)$, respectively, can be described by the summation of the factor stability bounds as:
$\|\varphi(u,t)-\tilde{\varphi}(u,t)\|=\sum_{p=1}^{P}{\mathcal{O}(\varepsilon_{p})}.$ (11)
@@ 3.3 Over-smoothing Analysis
$\text{where}\quad E(\mathbf{U})\vcentcolon=\frac{1}{2}\sum_{(i,j)\in\mathcal{E}}{\left\|\frac{\mathbf{u}_{i}}{\sqrt{\text{deg}_{i}}}-\frac{\mathbf{u}_{j}}{\sqrt{\text{deg}_{j}}}\right\|^{2}}=\operatorname{\operatorname{tr}}(\mathbf{U}^{\top}\hat{\mathbf{L}}\mathbf{U}).$ (13)
Inspired by (13), we extend the definition of Dirichlet energy to the tensorial case as follows:
Definition 3.8.
(Tensorial Dirichlet energy). By considering the normalized factor Laplacians $\{\hat{\mathbf{L}}_{p}\}_{p=1}^{P}$, we define the Tensorial Dirichlet energy for a tensor $\underline{\mathbf{U}}\in\mathbb{R}^{N_{1}\times\ldots\times N_{P}\times F}$ as:
$E(\underline{\mathbf{U}})\vcentcolon=\frac{1}{P}\sum_{f=1}^{F}{\sum_{p=1}^{P}{\operatorname{\operatorname{tr}}({\tilde{\mathbf{U}}^{\top}_{f_{(p)}}}{\hat{\mathbf{L}}_{p}}{\tilde{\mathbf{U}}_{f_{(p)}}})}},$ (14)
where ${\tilde{\mathbf{U}}_{f_{(p)}}}=\underline{\mathbf{U}}[:,\ldots,:,f]_{(p)}\in\mathbb{R}^{N_{p}\times\prod_{i\neq p}^{P}{N_{i}}}$, i.e., the $p$-th mode matricization of the $f$-th slice of $\underline{\mathbf{U}}$ in its $(P+1)$-th dimension.
Based on the previous research in the GSP literature [27, 33, 34], it follows that $E(\underline{\mathbf{U}})$ in (14) can be rewritten as $E(\underline{\mathbf{U}})=\operatorname{\operatorname{tr}}(\tilde{\mathbf{U}}^{\top}\hat{\mathbf{L}}\tilde{\mathbf{U}})$, where $\tilde{\mathbf{U}}\in\mathbb{R}^{[\Pi_{p=1}^{P}{N_{p}}]\times F}$, $\tilde{\mathbf{U}}[:,f]=\operatorname{\operatorname{vec}}(\underline{\mathbf{U}}[:,\ldots,:,f])$, and $\hat{\mathbf{L}}\vcentcolon=\frac{1}{P}\oplus_{p=1}^{P}{\hat{\mathbf{L}}_{p}}=\oplus_{p=1}^{P}{\left(\frac{\hat{\mathbf{L}}_{p}}{P}\right)}$. The next lemma shows interesting and applicable properties of $\hat{\mathbf{L}}$.
@@
Consider $P$ factor graphs with normalized adjacencies $\{\hat{\mathbf{A}}_{p}\}_{p=1}^{P}$ and Laplacians $\{\hat{\mathbf{L}}_{p}\}_{p=1}^{P}$. By constructing the product adjacency as $\hat{\mathbf{A}}\vcentcolon=\frac{1}{P}\oplus_{p=1}^{P}{\hat{\mathbf{A}}_{p}}=\oplus_{p=1}^{P}{\left(\frac{\hat{\mathbf{A}}_{p}}{P}\right)}$, a similar Cartesian form for the product Laplacian $\hat{\mathbf{L}}:=\mathbf{I}-\hat{\mathbf{A}}$ (with $\mathbf{I}\in\mathbb{R}^{[\Pi_{p=1}^{P}{N_{p}}]\times[\Pi_{p=1}^{P}{N_{p}}]}$ being the identity matrix) also holds as the following:
$\hat{\mathbf{L}}=\frac{1}{P}\oplus_{p=1}^{P}{\hat{\mathbf{L}}_{p}}=\oplus_{p=1}^{P}{\left(\frac{\hat{\mathbf{L}}_{p}}{P}\right)}.$ (15)
Besides, the spectrum of $\hat{\mathbf{L}}$ is spanned across the interval of $[0,2]$.
We observe in (14) that the kernel separability of Cartesian heat kernels also leads to the separability of analyzing over-smoothing in different modes of the data at hand. This is useful in cases of facing factor graphs with different characteristics. Next, we formally analyze over-smoothing in CITRUS. In our case, for the $l$-th layer with $H_{l}$ hidden MLP layers, non-linear activations $\sigma(\cdot)$ and learnable weight matrices $\{\mathbf{W}_{lh}\}_{h=1}^{H_{l}}$, we define:
@@
For the product Laplacian $\hat{\mathbf{L}}:=\frac{1}{P}\oplus_{p=1}^{P}{\hat{\mathbf{L}}_{p}}$ with domain-specific receptive field $t^{(i)}$ for the $i$-th factor graph and the activations $\sigma(\cdot)$ in (16) being ReLU or Leaky ReLU, we have:
@@
Theorem 3.10 shows that the factor graph with the smallest non-zero eigenvalue (spectral gap) multiplied by its receptive field dominates the overall over-smoothing. The less the spectral gap, the less probability the factor graph is connected [14]. So, we can focus on the (much smaller) factor graphs instead of the resulting massive product graph. These theoretical findings are experimentally validated in Section 4.2.
@@ 4 Experimental Results
In this section, we experimentally validate the theoretical discussions regarding the stability and over-smoothing analysis and use CITRUS in real-world spatiotemporal forecasting tasks. Similarly, we conduct ablation studies regarding CITRUS in spatiotemporal forecasting. We compare CITRUS against several state-of-the-art methods in forecasting. The results on the case of more than two factor graphs and descriptions of the state-of-the-art are provided in Appendix B and C, respectively, alongside the additional theoretical and experimental discussions in Sections D-I. The implementation codes are available at https://github.com/ArefEinizade2/CITRUS.